ABSTRACT. The relationship between the lower semicontinuity of the metric projections P_{G} onto a finite-dimensional subspace G of L_{1}, the Lipschitz continuity of P_{G}, the existence of continuous selections for P_{G}, and uniform strong uniqueness of P_{G} are studied. In particular, the lower semicontinuity of P_{G}, the Lipschitz continuity of P_{G}, and the uniform strong uniqueness of P_{G} are all equivalent. If P_{G} is lower semicontinuous, then P_{G} has a Lipschitz continuous selection. Moreover, if G is onedimensional, P_{G} has a continuous selection if and only if it has a Lipschitz continuous selection.